OpenStudy (anonymous):

What is the value of _8C_6? (Points : 1) 28 48 2520 20,160

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

\[a\text{C}b = \left( \begin{array}{c} a \\ b \\ \end{array} \right) = \frac{a!}{b!(a-b)!}\] \[8C6 = \frac{8!}{6!(8-6)!} = \frac{8*7*6*5*4*3*2*1}{6*5*4*3*2*1*2*1} = \frac{8*7\cancel{*6*5*4*3*2*1}}{\cancel{6*5*4*3*2*1*}2*1} \]\[\qquad= \frac{8*7}{2*1} =\]